Joan reads 3 pages in 2 minutes. How long will it take her to read 5 1/2 pages? Using formula d=r*t
Answer by TimothyLamb(4379)

use proportions:
---
3 pages / 2 minutes = 5.5 pages / x minutes
3/2 = 5.5/x
(3/2)x = 5.5
x = 5.5/(3/2)
x = 5.5*2/3
x = 3.6666666666 minutes
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php